OpenStudy (anonymous):

Solve by substitution method: 4x + 3y = 13 x + y = 4

OpenStudy (anonymous):

ok hold i on im typing it right now

OpenStudy (anonymous):

ok First you want to solve for x and y.4x+3y=13x+y=4, i have to solve for for y on the right. so i subtract 13x and 4 which gives me y= -13x+4 now you have to plug this in the left equation. 4x+3(-13x+4)= solve and you get x=1/4 and plug that into the left equation. 4(1/4)+3y=-1/3. x=1/4 and y=-1/3